4. What is the pH of a 0.50 M solution of HC2H3O2?
Answer #1

HC2H3O2 is acetic acid, usually written CH3COOH; it dissociates as follows:
CH3COOH ⇔ CH3OO- + H+, Ka = 1.8*10-5

0.50-x x x

If x moles of CH3COOH dissociate, x moles of H+ and x moles of CH3COO- are produced and the concentration of CH3COOH is reduced by x:

so Ka = x²/(0.50 - x)= 1.8*10-5

x²=1.8*10-5*(0.50 - x)

x will be small compared to 0.50, so Ka ≈ x²/0.50; x = √[0.50*1.8*10^-5] = 0.003

[H+] = x = 0.003;

pH =-log([H+]) =-log(0.003)=2.52

thank you